# The physical basis of teleological evolutions

### v4

@@ 4 The time-symmetric quantum causal loops – formalization
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary external observer. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her information about the solution of the problem.
(3) Complete the ordinary quantum description of the quantum algorithm by time-symmetrizing it. Mathematically, the quantum description of the process between the two correlated measurement outcomes remains unaltered. However, the description of the behavior of causality along it changes: the single causal process of the ordinary quantum description is replaced by a quantum superposition of time-symmetric causal loops.
@@
(4) Relativize the quantum description of the causal loops generated at point (3) with respect to the problem-solver. This produces the quantum description of the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
@@
4.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
@@
The usual quantum description of quantum algorithms is limited to the process of solving the problem, where the quantum speedup resides, and lacks the initial measurement. Of course, it is trivially incomplete: the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. We extend the usual quantum description to the upstream process of setting the problem. An initial measurement of the problem-setting, performed in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random. This is followed by the unitary evolution that computes the corresponding solution and the final measurement of the solution.
@@
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow444 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operantion that changes nothing for us. .
It is now Alice’s turn. The unitary part of her problem-solving action, $\hat{\mathbf{U}}_{1,2}$, produces the solution of the problem, of course, never changing the problem-setting. Note that, of $\hat{\mathbf{U}}_{1,2}$, it is sufficient to know that it does this, whereas we will never need to know how it does it. Of course, we must also know that there exists a unitary transformation that does this, Grover algorithm in the present case.
@@
Eventually, by measuring $\hat{A}$ (the content of $A$), which is already in one of its eigenstates, Alice acquires the solution with probability $1$ – the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement. Correspondingly, there is a unitary transformation between the initial and final measurement outcomes; the process between them is reversible despite comprising the final measurement.
4.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement555 More in general, one should assume that the problem-solver has a complete knowledge of the problem, thus also of the relation between problem-setting and solution. Knowledge of the problem-setting selected by the initial measurement would tell her also the solution. .
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by Bob and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following description of the quantum algorithm with respect to Alice:
@@
Note that this relativization relies on relational quantum mechanics $\left[33,34\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
4.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
@@
Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of the diagram.
@@
Summing up, mathematically, time-symmetrization leaves the unitary process between the two measurement outcomes unaltered but changes the structure of causality along it. The causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
4.4 Relativizing the causal loops with respect to the problem-solver
@@
Note that the entire causal loop is the zigzag diagram of table IV. It is a causal loop that changes its own past. The initial state in the top-left part of the diagram, where the state of register $B$ tells that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$. Since what happens along the time-symmetric causal loop obeys classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way, namely just one.
@@
The generalization of the above is straightforward. Note that, in all our considerations, we never needed to know the unitary transformation between the two measurement outcomes $\hat{\mathbf{U}}_{1,2}$. It sufficed to know that it unitarily changes its input state, where the two observables ($\hat{B}$ and $\hat{A}$ ) are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
First, the causal loops explanation of the quantum speedup, and the fact that the latter is quadratic in kind, remain there if we increase the number of quantum bits in the two registers of an even number and in an equal way. In fact, all our considerations remain identical. This means, in particular, Grover algorithm with number of drawers any even power of $2$.
The same, of course, holds, for any unitary evolution that builds up one-to-one correlation between two initially unrelated observables without changing the eigenvalue of the observable selected by the initial measurement. This will apply to the teleological evolutions addressed in the next section.
5 The physical basis of the notion of teleological evolution
We have seen, by the way rather simply, that the complete quantum description of quantum algorithms is a quantum superposition of time-symmetric causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in the optimal logical way benefiting of the advanced knowledge in question. In equivalent terms, the computational complexity of the problem to be solved or the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces. As already observed, in quantum search algorithms, also a random quantum search for the solution is endowed with (a suboptimal) quantum speedup.
Therefore, we can say that quantum algorithms are evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future. Of course, this is necessarily a causal loop and responds to the notion of teleological evolution. As is well known, before the scientific revolution, the evolutions of the living were often considered to be teleological in character. The teleological notion was then dismissed with the advent of modern science for the alleged absence of a physical basis. That was, of course, under the causal view of evolutions that is dominant in science also today.
Now, the fact that quantum algorithms turn out to be teleological evolutions should already cast a doubt on this way of seeing. It also, naturally, raises the question whether the time-symmetric quantum causal loops that stand at the basis of their teleological character could be the missing physical basis of the teleological character of natural evolutions.
@@ 5 The physical basis of the notion of teleological evolution
As already seen in the Rationale, under a quantum cosmological interpretation of the Anthropic Principle, the search for the value of life by the evolution of the universe under the fundamental physical laws and constants has the same form a quantum search algorithm. This search can therefore occur with quantum speedup and does even if it is random. Consequently, the evolution of a universe toward the value of life turns out to be teleological in character. Of course, in the universe we are in, what turns out to be teleological in character is the evolution of the universe toward sentient life.
@@
Eventually, as we have seen, the evolution of life on earth, in principle obtainable from that of the universe by tracing out all the rest, even if decohered, should inherit the quantum speedup and the consequent teleological character of the evolution of the universe toward life.
In the end, the notion that the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithm are also the physical basis of the teleological character of natural evolutions would seem to be physically plausible, this, of course, within the quantum cosmological view.
This section is a discussion of the many impacts that the present teleological view of natural evolutions might have on different areas of science and knowledge; of course, not forgetting the natural philosophy that gave rise to the notion of teleological evolution.
5.1 Decohered evolutions of life endowed with quantum speedup: how would they look?
All natural evolutions are necessarily described in a probabilistic way. We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
Instead, a decohered process endowed with quantum speedup must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection must make, in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the process would be difficult to distinguish from a classical stochastic process. Since also in this case the time-symmetric causal loops we are dealing with should be involved, let us call time-symmetric also the just outlined kind of stochastic evolution.
We mention what could be an emblematic example of time-symmetric stochastic evolution. It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
The usual explanation is of course a causal one: at first, feathers appeared through random mutations and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
The above is puzzling, of course. We should keep in our mind that we are under the quantum cosmological assumption that might make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as strange as the microscopic quantum world.
5.2 Teleological evolution of the universe toward life – a toy model
@@
A toy model of a teleological evolution of the universe toward life can be trivially mapped on the simplest instance of Grover algorithm.
@@
Initially, $\hat{B}$ should be in a quantum superposition (or, indifferently, mixture) of all its eigenstates (i.e of all the possible universes each with its fundamental physical constants). We can assume that also $\hat{A}$, the value of life observable, is initially in a quantum superposition of all its eigenstates.
@@
5.3 How to model a teleological natural evolution
@@
We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth after the beginning of the liquid water stage of it. The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. Of course, in the present case, this holds only in principle. However, it would be interesting to apply the canonical way to a toy model of the situation we are dealing with.
@@
One should identify the purely quantum evolution of a system (the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (the Earth). Then, one should obtain a decohered evolution of the subsystem from that of the system by tracing out all the rest. Note that the interaction of the subsystem with the rest of the system, in the present case, would do the opposite than decoherence. It would be what allows the evolution of the subsystem toward the goal to occur with quantum speedup as part of the purely quantum evolution of the system.
@@
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious (approximate) way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[35\right]$.
@@ 5.4 Impact on the natural sciences
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[36\right]$:
@@
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[37\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution. This in the sense that it better answers empirical evidence – of course the first canon of science. The only reason biologists would be against teleology in spite of empirical evidence would come from another science. We mean the notion – current in today’s physics – that a teleological evolution is physically impossible.
Of course, the sheer notion that quantum computation is teleological in character may already put a strain on the position of today’s biology: actually, although in the quantum world, teleological evolutions are physically possible. The strain should increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward life should be teleological in character. It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
@@
Almost because, on the one side, in recent years a group of evolutionary biologists $\left[38\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[35\right]$.

### v6

@@ 3.2 Formalization
(2) This point is propaedeutic to point (4). It shows how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement, which would give her full information about the solution of the problem555 We must assume that the problem-solver knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
(3) Complete the ordinary quantum description of the quantum algorithm by time-symmetrizing it.
@@
(4) Relativize with respect to the problem-solver the time-symmetric causal loops generated by the time-symmetrization of point (3). This produces the quantum superposition of causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
@@
3.2.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
@@
The usual quantum description of quantum algorithms is limited to the process of solving the problem, where the quantum speedup resides, and lacks the initial measurement. Of course, it is trivially incomplete: the canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. We extend the usual quantum description to the upstream process of setting the problem. An initial measurement of the problem-setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random. This is followed by the unitary evolution that computes the corresponding solution and the final measurement of the solution.
@@
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow666 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing. .
It is now Alice’s turn. The unitary part of her problem-solving action, $\hat{\mathbf{U}}_{1,2}$, produces the solution of the problem, of course, never changing the problem-setting. Note that, of $\hat{\mathbf{U}}_{1,2}$, it is sufficient to know that it does this, whereas we will never need to know how it does it. Of course, we must also know that there exists a unitary evolution that does this, it is the unitary part of Grover algorithm in the present case. Correspondingly, throughout this section, the unitary part of Grover algorithm will always remain just the symbol $\hat{\mathbf{U}}_{1,2}$.
@@
Eventually, by measuring $\hat{A}$ (the content of $A$), which is already in one of its eigenstates, Alice acquires the solution with probability $1$ – the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement. Correspondingly, there is a unitary transformation that connects the initial and the final measurement outcomes; the process between them is reversible despite comprising the final measurement (which does not change the quantum state).
3.2.2 Relativizing the ordinary quantum description of the quantum algorithm with respect to the problem-solver
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, namely an observer external to the quantum process. It is not a description for the problem-solver (Alice). In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the number of the drawer with the ball selected by Bob, which is also the solution of the problem – see the state of register $B$ in the bottom-left corner of Table I diagram. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement.
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by Bob and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following quantum description of the quantum algorithm with respect to Alice:
@@
Note that this relativization relies on relational quantum mechanics $\left[35,36\right]$, where the quantum state is relative to the observer. In the following, we will continue to call the quantum description with respect to the external observer the ordinary quantum description of the quantum algorithm. Instead we will always specify that the other quantum description is that with respect to the problem-solver.
3.2.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
@@
Also note that causality performs a causal loop that changes its own past; the outcome of its initial measurement of $\hat{B}_{l}$ is changed into the outcome of the initial measurement of $\hat{B}$ of the ordinary quantum description; compare the state of register $B$ under the left vertical arrow with the state in the bottom-left corner of Table III diagram.
@@
Summing up, mathematically, time-symmetrization leaves the quantum description of the unitary process between the two measurement outcomes unaltered but changes the structure of causality along the process. The causal process of the ordinary quantum description is replaced by a quantum superposition of causal loops.
3.2.4 Relativizing the causal loops with respect to the problem-solver
@@
The entire causal loop is the zigzag diagram of table IV. It is a causal loop that changes its own past. The initial state in the top-left part of the diagram, where the state of register $B$ tells us that Alice is completely ignorant of the number of the drawer with the ball, changes into the initial state in the bottom-left part of the diagram, where the state of register $B$ tells us that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$. Since what happens along the time-symmetric causal loop obeys classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way, namely just one.
@@
The generalization of all the above is straightforward. Of course, everything immediately generalizes to any number of drawers. Note that, in all our considerations, we never needed to know the unitary transformation ($\hat{\mathbf{U}}_{1,2}$) between the two measurement outcomes. It sufficed to know that there exists a unitary transformation that changes its input state, where the two observables $\hat{B}$ and $\hat{A}$ are unrelated, into its output state, where they are one-to-one correlated, without changing the eigenvalue of the observable selected by the initial measurement.
4 The possible physical basis of natural teleological evolutions
@@
We have seen in the Introduction how, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward the value of life (in particular, either sentient life or absence of life) under the fundamental physical laws and constants has the form of a quantum search algorithm. As such, it should be endowed with quantum speedup and would therefore be teleological in character even in the case that the search for the value of life is assumed to be random.
@@
In the universe we are in, the search for the value of life becomes search for life, of course endowed with the same quantum speedup. We have also seen that, as a consequence, the evolution of life on Earth, obtainable from that of the universe by tracing out all the rest, even if decohered, would inherit the quantum speed up of the evolution of the universe toward life and would therefore be teleological too. The same would necessarily hold for the evolutions of species, etc..
In this section, after encoding a toy model of the evolution of the universe toward the value of life on Grover quantum search algorithm, we will discuss the many impacts that the present physical notion of teleological natural evolution should have on different areas of science and knowledge.
4.1 Teleological evolution of the universe toward the value of life – a toy model
@@
A toy model of a teleological evolution of the universe toward the value of life can be trivially mapped on the simplest instance of Grover algorithm.
@@
Initially, $\hat{B}$ should be in a quantum superposition (or, indifferently, mixture) of all its eigenstates (i.e of all the possible universes each with its fundamental physical constants). We can assume that also $\hat{A}$, the value of life observable, is initially in an independent quantum superposition of all its eigenstates.
@@
We should note that the present toy model is less crude than it may seem. In fact, it immediately generalizes to any number $N$ of drawers. This means that any amount of information can be encoded in the process we are dealing with. Moreover, the unitary transformation $\hat{\mathbf{U}}_{1,2}$ between the initial and final one-to-one correlated measurement outcomes is any unitary transformation that does not change the eigenvalue selected by the initial measurement. Therefore, in principle, this toy model could even become a faithful discrete model of the actual evolution of the universe toward the value of life.
4.2 Decohered teleological evolutions: what would they look like?
The supposedly teleological evolutions of life would be, so to speak, around us in the macroscopic world. Decohered and endowed with quantum speedup it is natural to wonder what they would look like. Now, all natural evolutions are necessarily described in a probabilistic way. We can think of the stochastic description of a Darwinian evolution: an alternation of random changes in the species DNA pool each followed by natural selection of the DNA changes more suited to the environment. Of course, in a classical stochastic process, transition probabilities can depend on the past of the process but not on the future of it.
Instead, a decohered process teleological in character must have an attractor in its future. This means that those DNA mutations that in the future will survive natural selection because more fit to the environment must make, when they occur in the present, the transition toward them more probable. By the way, since the consequent teleological character of the evolution would be hidden in transition probabilities, the evolution would be difficult to distinguish from a classical stochastic evolution. Let us call time-symmetric the present teleological kind of stochastic evolution, given that its description would be generated by a time-symmetrization procedure.
Let us mention what could be an emblematic example of time-symmetric stochastic evolution. It is known that, for many millions of years, the ancestors of the flying birds were feathered dinosaurs who did not fly. Very light, strong but flexible, cave, and air-proof, feathers are of course ideal for flight. So, why anticipating something designed for flight millions of years in advance?
The usual explanation is of course a causal one: at first, feathers appeared by random mutation and survived natural selection as good thermal insulators, later they proved useful for speeding escape from predators, perhaps for gliding from trees to the ground, and only eventually they turned out to be what was needed to conquer the air.
The above explanation is puzzling, of course. We should keep in our mind that we are under the quantum cosmological assumption that should make macroscopic evolutions, the partial trace over all the rest of a quantum evolution of the universe, as puzzling as the microscopic quantum world.
4.3 How to model teleological natural evolutions
@@
We address the problem of how to model a decohered evolution endowed with quantum speedup – say that of the living on Earth after the beginning of the liquid water stage of it. The canonical way would be to obtain this evolution from that of the universe toward life by tracing out all the rest. Although, in the present case, this obviously holds only in principle, it would be interesting to apply it to a toy model of the situation.
@@
One should identify the purely quantum evolution of a system (standing for the universe) endowed with quantum speedup, where the achievement of the goal is mostly localized in a subsystem thereof (standing for the Earth). Then, one should obtain a decohered evolution of the subsystem from that of the system by tracing out all the rest. Note that the interaction of the subsystem with the rest of the system, in the present case, would do the opposite than decoherence. It would be what allows the evolution of the subsystem toward the goal to occur with quantum speedup as part of the purely quantum evolution of the system.
@@
Let us note that, since the liquid water stage of it, the evolution of life on Earth should have been almost completely independent from the evolution of the rest of the universe. If it were completely independent, it would be factorizable out of the evolution of the universe, thus itself quantum and endowed with quantum speedup. Therefore, an obvious, approximate, way of modeling the decohered evolution of the living on Earth endowed with quantum speedup is just to ignore the fact it is decohered and model it in a quantum way. There are encouraging precedents to this. For example, it has been shown that modeling in a quantum way the interplay between species of predators competing for the same preys allows to explain the empirically ascertained species’ asymptotic behavior in a way otherwise difficult to explain $\left[37\right]$.
@@ 4.4 Impact on the natural sciences
This tension is captured by the following sentence by the evolutionary biologist Haldane $\left[38\right]$:
@@
No biologist, of course, does not mean no scientist. A paradigmatic example is the American philosopher Nagel $\left[39\right]$, who argues that the principles that explain the emergence of life should be teleological, rather than materialistic or mechanistic in character.
We would put together Nagel’s view in support of the teleological character of the evolutions of the living and Haldane’s assertion that a biologist cannot live without teleology to affirm that the notion of teleological natural evolution is more “scientific” than the mainstream one of causal natural evolution. This in the sense that it better answers empirical evidence – of course the first canon of science. The only reason biologists would be against teleology despite empirical evidence would come from another science. We mean the notion – current in today’s physics – that a teleological evolution is physically impossible.
Of course, the sheer notion that quantum computation is teleological in character may already put a strain on the position of today’s biology: actually, although in the quantum world, teleological evolutions are physically possible. The strain might increase with the notion that, under a quantum cosmological interpretation of the Anthropic Principle, the evolution of the universe toward the value of life, toward life in the universe we are in, should be teleological in character. It should further increase with the consequent notion that also the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, even if decohered, should inherit the teleological character of the universe’s evolution.
@@
Almost because, on the one side, in recent years a group of evolutionary biologists $\left[40\right]$ expressed the need to rethink Darwin’s theory of evolution. That was after the discovery that some species’ evolution steps occurred far too quickly to be solely ascribed to random changes in the species’ DNA pool. Of course, the present teleological view of evolutions, with the associated quantum speedup, would offer a way out.
On the other side, there are scientists who have began to apply essentially quantum mechanical features like quantum entanglement and nonlocality to the macroscopic game of species, showing that this can explain some empirically ascertained features of it otherwise difficult to explain $\left[37\right]$.